### v3

# Diffusion Graph Neural Networks for Robustness in Olfaction Sensors and Datasets in Security Robotics

Abstract
Robotic odour source localization (OSL) is a critical capability for autonomous systems operating in complex environments. However, current OSL methods often suffer from ambiguities, particularly when robots misattribute odours to incorrect objects due to limitations in olfactory datasets and sensor resolutions. To address this challenge, we introduce a novel machine learning method using diffusion-based molecular generation to enhance odour localization accuracy that can be used by itself or with automated olfactory dataset construction pipelines. This generative process of our diffusion model expands the chemical space beyond the limitations of both current olfactory datasets and training methods, enabling the identification of potential odourant molecules not previously documented. The generated molecules can then be more accurately validated using advanced olfactory sensors, enabling them to detect more compounds and inform better hardware design. By integrating visual analysis, language processing, and molecular generation, our framework enhances the ability of olfaction-vision models on robots to accurately associate odours with their correct sources, thereby improving navigation and decision-making through better sensor selection for a target compound in critical applications such as explosives detection, narcotics screening, and search and rescue. Our methodology represents a foundational advancement in the field of artificial olfaction, offering a scalable solution to challenges posed by limited olfactory data and sensor ambiguities [^1] Code and data are made available to the community at the following URL: https://github.com/KordelFranceTech/OlfactionVisionLanguage-Dataset .
I Introduction
The human sense of smell is a complex and nuanced sensory modality, capable of distinguishing an extensive array of odourants. In recent years, artificial olfactory systems have been developed to detect sources of explosives, illegal drugs, and even perform forensic studies. These olfactory systems, such as electronic noses, utilize sensor arrays and pattern recognition algorithms to detect and classify volatile compounds [1, 2, 3, 4].
Despite sensor advancements, the development of machine learning models for olfactory perception faces significant challenges, primarily due to the scarcity of comprehensive olfactory datasets. Humans describe chemical odourants with lingual descriptors such as fruity and floral, but sensors interpret odourants by their direct chemical reactions, leading to a disparity in data training methods for olfactory sensors (e.g. how does one describe to an olfactory sensor the scent of explosives?). Datasets like those from Leffingwell [5] and GoodScents [6] have been instrumental in providing foundational results. However, due to the complexity in obtaining human-labeled odour descriptors over compounds, these datasets are limited in scope, lacking complete coverage of the vast chemical space associated with olfactory stimuli [7] and the necessary human controls to ensure objective measurements such as compensation for genetic variation in olfactory mechanisms [8], environmental context [9], health status [10, 11], age [12], and even potential neurocognitive conditions [13].
To address this data paucity, we propose an innovative diffusion-based graph neural network that allows one to minimize uncertainty the construction of olfactory datasets, particularly olfaction-vision datasets. Our motivation is driven by the desire to teach robots to navigate by scent and triangulate the source of explosives, building on methods from [14, 15, 16, 17]. Our research provides four primary contributions:
  1. 1.
    Our model aids in the selection of olfactory sensors for a desired target compound.
  2. 2.
    Our model increases the robustness of constructing olfactory datasets and we provide an open dataset and training method for use by the community.
  3. 3.
    In result of (1) and (2), our model reduces overall uncertainty and bias in olfactory navigation tasks.
  4. 4.
    As a result of (3), our methods directly increase accuracy and safety in olfaction-based security applications.
We combine visual analysis, language processing, and molecular generation to enhance the field of olfactory machine learning through the construction of robust olfaction datasets. By addressing the limitations of existing datasets and incorporating advanced sensor validation, we aim to advance the capabilities of artificial olfactory systems in accurately navigating by scent in complex environments.
II Background
The intersection of machine learning and olfaction presents a compelling frontier in artificial intelligence research. The human olfactory system’s ability to discern a vast array of scents is a complex process, involving intricate interactions between odourant molecules and olfactory receptors. Replicating this capability computationally necessitates sophisticated models and comprehensive datasets.
A significant hurdle in developing machine learning models for olfaction is the scarcity of extensive and high-quality datasets. Existing resources, such as the Leffingwell PMP 2001 and GoodScents databases, provide valuable information on various chemical compounds and their associated odour descriptors. However, these datasets are limited in scope and do not encompass the full spectrum of olfactory stimuli. This limitation hampers the training of robust machine learning models capable of generalizing across diverse olfactory inputs.
To address the data scarcity issue, researchers have explored the use of vision-language models (VLMs) to generate odour descriptors from images. VLMs, trained on large-scale image-text pairs, can infer contextual information from visual inputs and generate corresponding textual descriptions [18] [19]. By applying VLMs to images, it is possible to extract descriptive terms that may correlate with olfactory characteristics. However, the reliability of these descriptors is contingent upon the VLM’s exposure to olfactory-related data during training, the degree of which is not known. Consequently, the generated descriptors may lack specificity or accuracy in representing actual odours.
While efforts to construct olfaction-vision-language models (OVLMs) exist [20, 21], their underlying training data and training processes are not well standardized. By analyzing visual content in a standard computer vision dataset like, for example, the COCO dataset [22], VLMs can infer potential olfactory characteristics associated with depicted scenes or objects. For example, if a motorcycle is identified in an image, the VLM will reason that carbon monoxide is likely a present molecule and will list the aromas that describe carbon monoxide given the appropriate prompt. These inferred descriptors serve as a bridge to map visual information to chemical compounds, facilitating the identification of molecules likely present in a given image.
However, the reliance on VLMs introduces its own set of uncertainties, particularly regarding the accuracy and completeness of the generated odour descriptors. It is difficult to know the exact scope of chemistry contained in the VLM training data. To mitigate these concerns, we integrate a diffusion model trained on existing olfactory datasets and designed to generate novel molecular structures that correspond to the inferred odour descriptors, effectively expanding the chemical space beyond the limitations of current datasets. By doing so, we aim to capture a broader spectrum of odourant molecules, including those not previously documented. This is helpful in tasks oriented around olfactory sensors. Sensors that rapidly sample the air only screen for a select set of molecules. If a robot is given the task to navigate toward a specific odour, it has to be equipped with olfactory sensors that can target the specific molecules attributed to that odour. By understanding the target molecule’s near neighbors on the odour continuum, one can ensure to integrate sensors that address these molecules as well. As a simplified example, imagine that an unmanned aerial vehicle (UAV) is instructed to find the source of nitric oxide ($NO$), sometimes an indicator for the presence of explosives. $NO$ can, in some environments, quickly oxidize to $NO_{2}$. If the robot takes substantial time to navigate to the source of $NO$, it could be analyzing the wrong compound by missing the detection window, and at some point, may want to transition to detecting $NO_{2}$. The olfactory sensors (for example, metal oxide sensors) required to detect both $NO$ and $NO_{2}$ are different, and it would be prudent for the engineer to query which near neighbor compounds to $NO$ could be relevant for detection so the UAV could be equipped with such sensors. This issue is exemplified in tasks from [14] [15] [16] [17] [23]. We can make such a query automatic in our dataset construction, such that a separate machine learning model is queried for the nearest neighbors of any compound identified by the VLM. This not only informs better sensor selection, but can also partially reconcile erroneous compounds produced by the VLM.
In machine olfaction, the Shape Theory of Olfaction suggests that molecules that exhibit similar structural traits ”smell” similarly [24] [25] [26] . Although there is some support against this theory [27] [28] [29] [30], our proposed method assumes this theory is true, and we construct a model that builds off its principles. For example, if a VLM indicates that the odour descriptors ”fruity” and ”floral” are associated with an object in an image, one can use the GoodScent or Leffingwell datasets to look up particular molecules that exude such aromas. However, since these datasets are not comprehensive of all possible molecules, one can use the proposed diffusion model to generate ”near neighbors” of molecules that are structurally similar to those found in the existing datasets.
Figure 1: An example of how diffusion helps in olfactory modeling. A VLM describes aromas present in COCO images. Near neighbors of those aromas and molecules are then synthesized with our diffusion model for more comprehensive olfactory analysis of the visual scene.
To bridge the gap between odour descriptors and chemical compounds, diffusion models have emerged as a promising solution. Diffusion models are generative frameworks that learn to produce data samples by iteratively denoising random noise, guided by learned data distributions. In the context of molecular generation, diffusion models can be conditioned on textual inputs, such as odour descriptors, to generate novel molecular structures that correspond to the specified olfactory characteristics. This approach enables the exploration of chemical spaces beyond existing datasets, facilitating the discovery of new compounds with desired scent profiles. Recent studies have demonstrated the efficacy of text-guided diffusion models in generating molecules with specific properties, highlighting their potential in olfactory research [31] [32]. Further research by Lee, et al. [33] and Sisson, et al. [34] have demonstrated feasibility of other generative models in producing viable chemical compounds according to their desired odours or chemical bonds.
The validation of generated molecules necessitates empirical methods to confirm their olfactory properties. Advanced olfactory sensors offer a means to detect and analyze volatile compounds, providing data on their scent profiles. By comparing the sensor readings of generated molecules with the intended odour descriptors, researchers can assess the accuracy and relevance of the diffusion model’s outputs. In reality, this empirical validation is very difficult for reasons denoted in the Limitations section.
III Related Work
The integration of olfaction into robotic systems has garnered significant attention, particularly in the domain of robotic odour source localization (OSL). Traditional OSL approaches often rely solely on olfactory cues, which can be unreliable in complex environments due to factors like turbulent airflow and the presence of multiple odour sources. To address these challenges, recent studies have explored the fusion of multiple sensory modalities [35] [36].
For instance, a study introduced a fusion navigation algorithm that combines both vision and olfaction-based techniques. This hybrid approach addresses challenges such as turbulent airflow, which disrupts olfaction sensing, and physical obstacles inside the search area, which may impede vision detection. The methodology includes a custom-trained deep-learning model for visual target detection and a moth-inspired algorithm for olfaction-based navigation. Experimental results demonstrate that this vision and olfaction fusion algorithm significantly outperforms vision-only and olfaction-only methods, reducing average search time by 30% and 54%, respectively [35] [36].
In parallel, advancements in machine learning have facilitated the generation of novel molecules based on textual descriptions. Text-guided diffusion models, such as TextSMOG, have been developed to generate small molecules conditioned on textual inputs. These models enhance both stability and diversity in molecule generation, capturing and utilizing information from textual descriptions effectively [31].
Furthermore, the application of diffusion models in molecular design has been extensively reviewed, highlighting their potential in generating molecules with desired properties. These models operate by simulating the gradual degradation of a data distribution and learning its reverse process to generate new samples. Their success in visual domains has inspired researchers to explore their potential for molecular generative tasks, making them central to molecular design [32]. Work from Lee, et al. in [33] show how graph neural networks can be used to construct principal odour maps for deducing probable aromas for a given molecule. Our work here with graph neural networks strives to align with many of their assumptions, but leverages diffusion to overcome uncertainties associated with changing perceivable aromas among different concentrations of the same compound as exemplified by Longin, et al. [37].
Despite these advancements, challenges persist in accurately associating odours with their correct sources in complex environments. Our proposed methodology builds upon these existing works by integrating vision-language models to extract odour descriptors from images and employing diffusion models to generate corresponding molecular structures. This approach aims to enhance the robot’s ability to disambiguate odour sources, improving navigation and decision-making in environments where olfactory cues are essential.
IV Methodology
IV-A Diffusion Model Selection
Traditional methods for exploring chemical similarity, such as nearest neighbors using molecular fingerprints, are effective for retrieving known molecules from existing datasets. These methods are particularly well-suited for tasks that involve classification, clustering, or similarity-based search. However, they fall short when the goal is to generate entirely new molecules that possess specific olfactory properties — especially combinations of scent descriptors not commonly observed in known databases.
In contrast, diffusion models are generative frameworks capable of sampling novel molecular graphs from a learned data distribution. When trained on molecular structures annotated with multi-label scent descriptors, diffusion models can conditionally generate new candidate molecules that reflect specified olfactory profiles (e.g., floral, musky, woody). This goes beyond retrieval — it enables creation, offering several key advantages:
Unlike k-nearest neighbor methods, diffusion models do not rely on existing database entries. They have generative capability and can synthesize new, chemically valid molecular structures that conform to desired scent characteristics. Diffusion models also support multi-label conditioning, allowing fine control over the generated output. One can, for instance, request molecules that are simultaneously fruity, earthy, and sweet — even if no molecule with that combination exists in the training data. By modeling a distribution over molecular structures, diffusion models allow for sampling of diverse outputs that satisfy the same olfactory constraints, improving the breadth of chemical space explored during generation. The denoising process intrinsic to diffusion models implicitly learns smooth transitions in molecular structure and scent characteristics, enabling interpolation and optimization across scent manifolds. The diffusion framework can naturally be extended to incorporate additional modalities such as synthetic accessibility, toxicity, or volatility, making it suitable for multi-objective molecule design.
While diffusion models require more computational resources to train and sample compared to nearest neighbors, their generative power and flexibility make them particularly well-suited for olfactory molecule discovery and scent-driven innovation. These properties are essential for developing novel fragrance molecules, identifying rare scent combinations, and automating early-stage formulation in perfumery, flavor chemistry, and environmental sensing.
IV-B Graph Neural Networks
Our use of equivariant graph neural networks (EGNNs) enhances our diffusion architecture. EGNNs model atoms as both node features (e.g., atomic number), and continuous 3D positions. As a result, they learn from spatial relationships like distance and angle, and their outputs are equivariant to translation and rotation. In turn, this enables geometry-aware bond type inference, better generalization to shape-driven scent effects, and modeling conformational flexibility, important for subscribing to the Shape Theory of Olfaction.
V Model Architecture
V-A Diffusion Framework
Our generative framework is built on a conditional denoising diffusion model, enhanced by an EGNN architecture. This design is tailored to generate novel molecular structures in accordance with specific olfactory descriptors. Unlike traditional approaches that rely solely on molecular fingerprints or 2D-graph representations, our model incorporates 3D-molecular geometry and jointly learns both atomic identities and bond structures through time-reversible denoising steps.
Each molecule is represented as a graph $G=(V,E)$, where nodes correspond to atoms and edges represent chemical bonds. We enrich this representation with 3D coordinates $\mathbf{r}_{i}\in\mathbb{R}^{3}$ for each atom, computed using RDKit’s ETKDG embedding method [38]. Atoms are initially encoded as scalar atomic numbers, while bond types are annotated using categorical labels (single, double, triple, aromatic).
The model conditions generation on multi-label olfactory descriptors such as floral, musky, and fruity on a separate olfactory conditioner, which is a simple feedforward neural network. These descriptors are encoded as multi-hot binary vectors and projected into a continuous latent space via a learnable feedforward layer. This conditioning vector is then concatenated with both the node features and a time embedding to guide the diffusion process.
We adopt the standard forward diffusion process from a denoising diffusion probabilistic model (DDPM), where noise is incrementally added to the atomic node features over $T$ timesteps. For a given molecule, the clean node features $x_{0}$ are perturbed into a noisy version $x_{t}$ using:
$x_{t}=x_{0}+\sqrt{\beta_{t}}\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,I)$
where $\beta_{t}$ is the variance schedule and $t\in[1,T]$. In our implementation, we simplify this by scaling the noise linearly with $t$.
The training objective is to predict the noise $\epsilon$ added to $x_{0}$ given the noisy input $x_{t}$, the timestep $t$, and a conditioning vector representing the olfactory labels.
In other words, the model is trained to reverse this process by predicting the original noise component $\epsilon$ from the noisy input $x_{t}$, conditioned on both time $t$ and the olfactory label embedding. A simple linear time embedding module is employed to encode timestep information.
Molecules in the GoodScents dataset are annotated with multi-label binary vectors $y\in\{0,1\}^{L}$, where $L$ is the number of unique olfactory descriptors (e.g., floral, musky, fruity). These labels are projected into a latent conditioning space via a feedforward projection:
$\mathbf{c}=\text{Linear}(y)$
This conditioning vector $\mathbf{c}$ is concatenated with both the node features and timestep encoding before being processed by the denoising model.
The core denoising mechanism comprises two stacked EGNN layers. Each EGNN layer performs equivariant message passing by computing distances and directional vectors between atoms, then applies the learned neural network to both update node features and apply geometry-aware coordinate shifts. Formally, for a node pair $(i,j)$, the message $m_{ij}$ is computed using:
$m_{ij}=\text{MLP}_{\text{node}}\left([x_{i},x_{j},\|\mathbf{r}_{i}-\mathbf{r}_{j}\|]\right)$ (1)
A corresponding coordinate update is also computed:
$\Delta\mathbf{r}_{i}=\sum_{j\in\mathcal{N}(i)}\text{MLP}_{\text{coord}}(\|\mathbf{r}_{i}-\mathbf{r}_{j}\|)\cdot(\mathbf{r}_{i}-\mathbf{r}_{j})$ (2)
These updates preserve equivariance under rigid-body transformations, allowing the model to respect the geometric nature of molecular structures. The aggregated node messages are summed and added to the current feature representation $x_{i}$, while the position updates are added to $\mathbf{r}_{i}$.
In addition to denoising atomic identities, the model includes a parallel bond type predictor. For each edge in the graph, it concatenates the embeddings of the two endpoint nodes and predicts a bond type using a four-way softmax classifier. This classifier is supervised via a cross-entropy loss computed against ground truth bond labels.
The training objective loss $\mathcal{L}_{\text{total}}$ thus consists of two components: a mean squared error (MSE) loss $\mathcal{L}_{\text{MSE}}$ for node denoising, and a cross-entropy loss $\mathcal{L}_{\text{CE}}$ for bond classification:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{MSE}}(x_{t},\hat{x}_{0})+\mathcal{L}_{\text{CE}}$ (3)
Figure 2: Overview of the olfactory-guided diffusion model architecture. Atom features and 3D coordinates evolve over time via EGNN layers, while bond types are predicted in parallel. The model is conditioned on olfactory descriptor embeddings.
Once trained, generation begins from Gaussian noise. A noisy vector $x_{T}$ and random 3D coordinates are initialized and iteratively denoised over $T$ steps. After the final step, atomic predictions are rounded and filtered to chemically reasonable elements (e.g., C, N, O, S, Cl). The model also predicts bond types, and a molecular graph is assembled and sanitized using RDKit. If valid, the final structure is converted to a canonical SMILES string and visualized. Conditioning vectors can be customized (e.g., setting ‘floral‘ and ‘fruity‘ to 1, others to 0) to generate novel molecules with targeted scent profiles.
To enforce chemical validity, we include filtering rules during decoding. Atoms with unstable valence configurations or rare atomic numbers are discarded. Additionally, we apply nan_to_num to model outputs to prevent numerical instabilities. Molecules are validated and visualized using RDKit’s sanitization and rendering tools. During training, we apply temperature-based softmax to the bond logits $\sigma_{bond}$ according to the following:
$\sigma_{scaled}=\sigma_{bond}/\tau$ (4)
Where $\tau$ represents temperature. This helps reduce overly random predictions and improve structural validity. Figure 2 illustrates the architecture and generation process defined above.
In summary, our model provides a 3D-aware, scent-conditioned molecular generation system grounded in diffusion theory and equivariant geometry. It enables controlled exploration of molecular space according to sensory targets and opens a new pathway for olfactory-driven molecular design.
V-B Molecular Validation
Our diffusion model needs a series of checks and balances to ensure no nonsensical molecules are produced. To do this, we incorporate a series of validation steps.
Firstly, we validate the atomic number range. The periodic table only defines elements up to atomic number 118. Diffusion models might predict out-of-range values (e.g., negative or exaggerated atomic numbers) due to noise. Skipping such atoms ensures that the generated molecule remains within the bounds of known chemistry. This validation prevents the addition of chemically nonsensical atoms to the molecule.
We also ensure edge deduplication by avoiding the creation of multiple redundant bonds between the same pair of atoms. Molecular graphs are undirected and generally allow only one bond per atom pair (with varying types). This check ensures graph realism and prevents downstream errors in RDKit, which will reject duplicate edges unless explicitly defined as aromatic or resonance structures.
To validate the bond type, we add a heuristic to bond type inference. This validation converts continuous-valued output of the diffusion model into discrete bond types (single, double, triple). Since the model doesn’t explicitly predict bond types, this heuristic uses the difference in atomic number features as a proxy. While not chemically perfect, it introduces structural diversity and prevents over-simplified molecules (e.g., with only single bonds). Our model also reconciles bonding errors that may occur during inference. Even after edge deduplication and bond-type inference, some atom pairs may still be chemically incompatible for bonding (e.g., noble gases or already saturated atoms). This try/except block avoids crashes and skips invalid bond additions.
Other sanitization techniques of our model involve the addition of implicit hydrogens, validation of molecular valence and connectivity, and generation of correct kekulé or aromatic representations. Sanitization is critical for ensuring that the generated molecule is chemically plausible and renderable, especially for descriptor computation or visualization.
Even with the above rules in place, invalid molecules may still result from our diffusion model. Molecules that pass validation up to this point are further checked to assess unbalanced aromatic rings or invalid formal charges. This acts as a final safeguard before converting the final molecule to its SMILES and prevents flawed outputs from leaking into evaluation metrics or dataset augmentation.After SMILES conversion, we perform one final check to ensure human-readability, existence of the SMILES against a known database, and compatibility with downstream chemoinformatics tools.
VI Experiments & Results
Figure 3: Loss curves of the final trained models for both the constrained and unconstrained variants.
The first of our experiments was to train the EGNN and diffusion model on the LeffingWell and GoodScent datasets and evaluate the efficacy with the VLM. We borrow the dataset from [33] as it succinctly combines both datasets into a methodical format. The resulting dataset is nearly 5000 samples in size. While this is a small dataset in comparison to modern machine learning tasks in other modalities like image recognition, it is rare to have a dataset this size in the realm of olfaction. We use an 80-20 train-test split in which we evaluate our models over the indicated data. We find that diffusing between 800-1200 steps is ideal, and perform our experiments at 1000 steps. We train the EGNN for 1000 epochs with an embedding dimension of size 8. We train two models in the exact same manner with the exception that one model is unconstrained in that it can generate molecules from any available atom; the other model is ”constrained” in that it can only generate molecules from the following atoms: C, N, O, F, P, S, and Cl. The training loss for both models is shown in Figure 3. The final results are summarized in Table I.
We expected to see better performance by constraining the diffusion to a specific set of atoms. However, this actually resulted in worse performance, generating validated molecules over less than 10% of samples. We suspect this is because the permutation of such a small set of elements did not provide a diverse enough dataset from which the diffusion model could generate neighboring compounds in most cases. While the unconstrained model is more complex, we find that the increased sample space in combination with a longer training time yielded better performance than the constrained model and we leverage that model to communicate our final results.
Although molecules are generated for every set of descriptors that are input to the model, we find that only 27.71% of molecules diffused from our test set pass our molecular validation checks. This is to be expected and quite welcome as we suspect not every permutation of identified odour descriptors returns a large array of near-neighbor molecules. Note that our training data does not include all feasible compounds and their odour descriptors  [^2] For more information on the limitations of the Leffingwell and GoodScent datasets, please consult [5] and [6], respectively. We acknowledge that our approach here inherits any limitations noted in these datasets, including those attributed to the subjectivity of human-produced olfactory labels. However, we hope that our diffusion model partially reconciles some of the uncertainty associated with these olfactory labels..
We then test our proposed model with the VLM over the full test data of the COCO dataset [22]. We select GPT-4o [39] as our VLM and prompt it to caption each image with odour descriptors that are suspected to be present. An array of odour descriptors is then associated with each image in the test set and input into our diffusion-EGNN model. We obtain more permutations of odour descriptors than what are available in our training data, which leads to more possible molecules being produced than those experienced in our training set. Because of this, we note a 28.20% success rate in diffusing near-neighbor molecules from VLM odour descriptors after validation checks are passed. This marks a pleasing transfer of training accuracy to test accuracy and amounts evidence for us to suggest that the aromas for the sample of molecules provided in the consolidated Leffingwell and GoodScent datasets may be good proxies of the much larger molecular distribution.
TABLE I: Final results: Percentage of test samples with chemically-viable generated molecules.
Dataset Constrained Unconstrained
GS & LW Test Set [33] $10\%$ 27.71%
COCO Test Set [22] $10\%$ 28.20%
VII Limitations
While our integrated framework combining VLMs and diffusion-based molecular generation offers a novel approach to enhancing robotic odour source localization, it is essential to acknowledge its limitations and the challenges that persist in scent-based navigation. VLMs, though powerful in bridging visual and textual modalities, are not specifically trained on olfactory datasets. Consequently, their ability to generate accurate and comprehensive odour descriptors from images is constrained. This limitation can lead to incomplete or imprecise odour representations, affecting the grounding of subsequent molecular generation process and therefore the methods we delineate here. Moreover, VLMs may struggle with contextual reasoning and spatial understanding, which are crucial for accurately associating odours with their sources in complex environments. This can be analogously observed from the work of Xie, et al. in [40] where they attempt to infer sound from images using VLMs. For example, we noticed in our training that the VLM tends to associate carbon monoxide to the presence of a vehicle in the COCO image, but does not consider the fact that the vehicle may be electric.
The diffusion model, trained on existing olfactory datasets, aims to generate novel molecular structures corresponding to inferred odour descriptors. However, the quality and diversity of the training data directly influence the model’s generative capabilities. Given the limited scope of current olfactory datasets, the diffusion model may not capture the full spectrum of possible odourants or molecules, potentially leading to gaps in odour representation. This is exemplified in the fact that we observed more valid molecules generated with the VLM than with the training set.
Additionally, our method gives heavy credence to the Shape Theory of Olfaction. If this theory is proven untrue, it may invalidate the efficacy of our method. The generated molecules require empirical validation to confirm their olfactory properties which can be accomplished through various olfaction sensors. However, this can be very resource-intensive as the instrumentation required to validate the presence of compounds is expensive. In addition, obtaining all possible molecules over which to evaluate said sensors can be restrictive due to regulations and required licenses. Finally, even if one could obtain testing samples of all possible compounds in a unanimous quantity, it is not enough to test each compound individually. The combination and interaction of certain compounds produce entirely new odour descriptors which are not yet entirely predictable. Rapidly quantifying the presence of compounds within an air sample and the aromas attributed to them is a known problem within olfaction; this underscores the need for more datasets as proposed here.
Another significant challenge in robotic OSL is the accurate association of detected odours with their correct sources. Environmental factors such as airflow dynamics, presence of multiple odour sources, and obstacles can cause odour plumes to disperse unpredictably, leading to potential misattribution of odours to incorrect objects. While our framework enhances the robot’s ability to infer and generate potential odourant molecules, it does not entirely eliminate the possibility of such misassociations. Therefore, the system may still encounter difficulties in environments with complex odour landscapes.
Implementing the proposed framework in real-time robotic systems poses computational challenges. The integration of VLMs, diffusion models, and olfactory sensors requires efficient processing capabilities to ensure timely decision-making during navigation. Latency in processing can hinder the robot’s responsiveness, especially in dynamic environments where rapid adaptation is necessary. It should be noted that the proposed method is intended to aid in the selection of which olfactory sensors should be integrated on a robot to navigate to a particular compound, prior to deployment.
In summary of the above, we acknowledge that there are inherent limitations of our proposed methodology, but hope that it can be used to generate highly probable compounds for given aromas when constructing vision-olfactory datasets and informing sensor selection in olfactory robotics tasks.
VIII Conclusion
Machine olfaction is still a young area of AI and robotics that receives disproportionate attention and standardization in comparison to other modalities such as computer vision and audition. This creates several opportunities for dataset construction, standardization, and benchmarking. More attention is especially needed to adapt these methods to complex environments such as robotic OSL to toxic compounds. Our methodology represents a foundational step towards improving sensor selection in olfactory robotics by integrating visual, linguistic, and olfactory data. We note opportunities to utilize our method as both a standalone algorithm and as part of an automation pipeline used to construct olfactory datasets via VLMs. Our method enhances the robot’s ability to generate and associate odour descriptors with near-neighbor molecular structures, thereby aiding in more accurate odour localization. However, it is not a comprehensive solution. The limitations outlined above highlight the opportunities for continued research to address the complexities inherent in olfactory perception. This manuscript evaluates one method to reconcile known shortcomings in both the construction of olfaction-vision datasets and the selection of sensors for robotic OSL, but we hope it inspires more researchers in the field to optimize our solution and address the vast array of opportunities within artificial olfaction.